At the beginning of the current fiscal year, the balance sheet of Moondrop Co. showed liabilities of $760,000. During the year liabilities increased by $20,000, assets increased by $110,000, and paidin capital increased by $40,000 to $330,000. Dividends declared and paid during the year were
$120,000. At the end of the year, stockholders' equity totaled $804,000. Calculate net income or loss
for the year along with the other amounts.
Required: Calculate the value of the 7 numbers.

Assets = Liabilities + Paid In Capital + Retained Earnings
Beginning $(5) = $760,000 + (4) (6)
Changes +$110,000 +$20,000 +$40,000 (7)
Ending (3) (2) $330,000 (1)

1)

2)

3)

4)

5)

6)

7)
Answer #1

Assets=Liabilities+Paid in capital+Retained Earnings

Assets = Liabilities + Paid in Capital + Retained earnings
Beginning 1,474,000 = 760000 + 290,000 + 424,000
changes 110,000 = 20,000 + 40,000 + 50,000
Ending 1,584,000 = 780,000 + 330,000 + 474,000
  • Ending Liabilities=Opening Liabilities+ Change

=760,000+20,000

=780,000

  • Beginning Paid in Capital=Ending capital-Change

=330,000-40,000

=290,000

  • Retained earning=Total Stockholders' equity-Paid in capital

=804,000-330000

=474000

  • Ending assets=Total Libilities+Total Stockholders' equity

=780000+804000

=1,584,000

  • Beginning Assets=Ending Assets-Change

=1,584,000-110,000

=1,474,000

Net Income:-

Change In retained earnings =50,000

Retained earning is increased due to Net Income and decrease due to Dividend paid.

Dividend paid during the year=120,000

Net Income =50,000+120,000

=$170,000

"Please Rate if it helps you or let me know if anything else''